OpenStudy (anonymous):

find all solutions in the interval (0,2pi) 3sin(x-pi/5)=0

OpenStudy (aaronq):

\(3sin(x-\dfrac{\pi}{5})=0\) \(sin(x-\dfrac{\pi}{5})=\dfrac{0}{3}\) \(sin^{-1}(sin(x-\dfrac{\pi}{5}))=sin^{-1}(\dfrac{0}{3})\) \(x-\dfrac{\pi}{5}=sin^{-1}(\dfrac{0}{3})\) \(x=sin^{-1}(\dfrac{0}{3})+\dfrac{\pi}{5}\)